OpenStudy (anonymous):

sqrt(1-y^2)dx+sqrt(1-x^2)dy=0

OpenStudy (amistre64):

looks quite seperable to me ....

OpenStudy (anonymous):

how do we solve it though.. yup its separable but stuck up after the first step

OpenStudy (amistre64):

1/sqrt(1-a^2) maybe try some trig substitutions?

OpenStudy (anonymous):

tried ...i get the integrated form of the answer but not able to reach final conclusion

OpenStudy (amistre64):

hmm, let sin(u) = x, cos(u) du = dx ...sin(v) = y, cos(v) = dv sqrt(1-y^2)dx+sqrt(1-x^2)dy=0 cos^2(v)du+cos^2(u)dx sec^2(u)du= -sec^2(v) dv isnt tan the integral of sec^2? tan(u)= -tan(v) + C u= arctan[ -tan(v) + C ]

OpenStudy (amistre64):

arcsin(x)= arctan[ -tan(arcsin(y)) + C ] x= sin {arctan[ -tan(arcsin(y)) + C ] } might need to draw some pictures to look thru the trig parts

OpenStudy (anonymous):

no probs ..thanku :)

OpenStudy (amistre64):

youre welcome ... thats my idea at least :) good luck